Compact discs cost $16.99 each and cassette tapes cost $8.99 each. If the total bill equaled $78.93, how many of each were bought?
Answer by checkley77(12844)

16.99x+8.99y=78.93
7*99=#.93
Therefore:
x+y=7 or x=7-y
16.99(7-y)+8.99y=78.93
118.93-16.99y+8.99y=78.93
-8y=78.99-118.93
-8y=-40
y=40/8=5 tapes were bought.
x=7-5=2 discs were bought.
Proof:
16.99*2+8.99*5=78.93
33.98+44.95=78.93
78.93=78.93